OpenStudy (anonymous):

4 over 3x + 2 over 3 = 6

OpenStudy (anonymous):

4/3x + 2/3 = 6 First let's convert that 6 to a fraction with the same denominator as the rest. 4/3x + 2/3 = 18/3 Subtract 2/3 from both sides of the equation 4/3x = 16/3 Now multiply by the inverse of 4/3 on both sides to get x by itself 4/3x (3/4) = (16/3)(3/4) x = 48/12 = 4